OpenStudy (anonymous):

perform the operations and simplify: 3x+4/x^2-16 - 2/x-4

OpenStudy (pfenn1):

Is this your expression?\[\frac{3x+4}{x^2-2} - \frac{2}{x-4}\]

OpenStudy (anonymous):

how do i solve this problem?

OpenStudy (pfenn1):

I have to go eat dinner but maybe @dpaInc can help you.

OpenStudy (anonymous):

mmm. what are we having!!! pizza?

OpenStudy (anonymous):

ok.. hello?

OpenStudy (anonymous):

i'm here

OpenStudy (pfenn1):

What is the LCD?

OpenStudy (anonymous):

i don't know

OpenStudy (pfenn1):

well, let's try a LCD of\[(x^2-2)(x-4)\]What we get is\[\frac{3x+4}{x^2-2} - \frac{2}{x-4}=\frac{(3x+4)(x-4)-2(x^2-2)}{(x^2-2)(x-4)}\]

OpenStudy (pfenn1):

Can you multiply out the numerator?\[(3x+4)(x-4)-2(x^2-2)=\text { ?}\]

OpenStudy (anonymous):

if there was an equal sign rather than the subtraction sign between the two only then you can do that. That is why this question is puzzling. you are cross multiplying...is that a legal move?

OpenStudy (pfenn1):

@mellylol, am I cross multiplying? I thought I was coming up with a LCD. What should I have done?